OpenStudy (anonymous):

Write an equation for the line passing through B(0, 3) and C(–3, 4).

OpenStudy (anonymous):

U know two-point formula

OpenStudy (anonymous):

@samalami ???

OpenStudy (anonymous):

point slope formula? @sauravshakya

OpenStudy (anonymous):

Ok..... that will work too

OpenStudy (anonymous):

First find the slope

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

how would I find that without only points?

OpenStudy (anonymous):

Isnt coordinates of B and C are given

OpenStudy (anonymous):

then, slope=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

Did u get slope? @samalami

OpenStudy (anonymous):

slope is 1/-3?

OpenStudy (anonymous):

Yep. its -1/3

OpenStudy (anonymous):

Now, the required equation of line is (y-y1)=m (x-x1)

OpenStudy (anonymous):

Use any point for x1 and y1 and m=slope

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

then from there simplify the equation?

OpenStudy (anonymous):

YEP

OpenStudy (anonymous):

thank you <3 <3

OpenStudy (anonymous):

WELCOME